# Operator World Models for Reinforcement Learning

### v1

@@ B.1 Convergence rates of PMD
As usual, in this proof we denote the estimated and exact action-value functions as $q_{n}({P}_{t}):=\hat{q}_{\pi_{t}}=({Id}-\gamma{T}_{n}{P}_{t})^{-1}r_{n}$ and $q({P}_{t}):=q_{\pi_{t}}=({Id}-\gamma{T}{P}_{t})^{-1}r$, respectively. From hypothesis, Alg. 1 is well-defined since all policies it generates are $({\mathcal{G}},{\mathcal{F}})$-compatible. The resulting sequence of policies $(\pi_{t})_{t\in\mathbb{N}}$ are generated via the update rule 13 on the inexact action-value functions $q_{n}({P}_{t})$, as defined in 11. As the update rule 13 is a (measurable) minimizer of B.1 when $D$ equals the Kullback-Leibler divergence, the three-points Lemma B.2 with $\pi(x)=\pi_{*}(x)$ yields
$\left[({P}_{*}-{P}_{t+1})q_{n}({P}_{t})\right](x)\leq\frac{1}{\eta}D(\pi_{*}(x);\pi_{t}(x))-\frac{1}{\eta}D(\pi_{*}(x);\pi_{t+1}(x)).$
Adding and subtracting the term $\left[({P}_{*}-{P}_{t+1})q({P}_{t})\right](x)$, and bounding the remaining difference as $\left[({P}_{*}-{P}_{t+1})(q({P}_{t})-q_{n}({P}_{t}))\right](x)\leq 2\varepsilon_{t}$ – see the derivation of B.8 – one gets
$\left[({P}_{*}-{P}_{t+1})q({P}_{t})\right](x)\leq 2\varepsilon_{t}+\frac{1}{\eta}D(\pi_{*}(x);\pi_{t}(x))-\frac{1}{\eta}D(\pi_{*}(x);\pi_{t+1}(x)).$
Adding and subtracting ${P}_{t}q({P}_{t})$ on the left side gives
$\left[({P}_{*}-{P}_{t})q({P}_{t})\right](x)\leq\left[({P}_{t+1}-{P}_{t})q({P}_{t})\right](x)+2\varepsilon_{t}+\frac{1}{\eta}D(\pi_{*}(x);\pi_{t}(x))-\frac{1}{\eta}D(\pi_{*}(x);\pi_{t+1}(x)),$
and integrating with respect to the positive measure $({Id}-\gamma{P}_{*}{T})^{-*}\nu$ and using the performance difference Lemma A.4 on the left hand side one has
$\begin{split}J({P}_{*})-J({P}_{t})\leq&\frac{1}{1-\gamma}\left\langle{({P}_{t+1}-{P}_{t})q({P}_{t})+2\varepsilon_{t}},{d_{\nu}({P}_{*})}\right\rangle\\ &\frac{1}{\eta(1-\gamma)}\left\langle{D(\pi_{*};\pi_{t})-D(\pi_{*};\pi_{t+1})},{d_{\nu}({P}_{*})}\right\rangle,\end{split}$ (B.9)
where we used A.8 on the right-hand-side terms. Since $({P}_{t+1}-{P}_{t})q({P}_{t})+2\varepsilon_{t}\geq 0$ because of B.7, we can use fact (1) from Lemma A.6 with $({Id}-\gamma{P}_{t+1}{T})^{-1}$ and the performance difference Lemma A.4 to get
$\begin{split}\left\langle{({P}_{t+1}-{P}_{t})q({P}_{t})+2\varepsilon_{t}},{d_{\nu}({P}_{*})}\right\rangle&\leq\left\langle{({Id}-\gamma{P}_{t+1}{T})^{-1}\left[({P}_{t+1}-{P}_{t})q({P}_{t})+2\varepsilon_{t}\right]},{d_{\nu}({P}_{*})}\right\rangle\\ &=\left\langle{{P}_{t+1}q({P}_{t+1})},{d_{\nu}({P}_{*})}\right\rangle-\left\langle{{P}_{t}q({P}_{t})},{d_{\nu}({P}_{*})}\right\rangle+\frac{2\varepsilon_{t}}{1-\gamma}.\\ \end{split}$
@@
$\begin{split}\sum_{t=0}^{T-1}J({P}_{*})-J({P}_{t})\leq&\frac{1}{1-\gamma}\left(\left\langle{{P}_{T}q({P}_{T})},{d_{\nu}({P}_{*})}\right\rangle-\left\langle{{P}_{0}q({P}_{0})},{d_{\nu}({P}_{*})}\right\rangle\right)+\frac{2}{(1-\gamma)^{2}}\sum_{t=0}^{T-1}\varepsilon_{t}\\ &\frac{1}{\eta(1-\gamma)}\left\langle{D(\pi_{*};\pi_{0})-D(\pi_{*};\pi_{T})},{d_{\nu}({P}_{*})}\right\rangle.\end{split}$
Using facts (3) and (4) from Lemma A.6 we have that the terms $\left\langle{{P}q({P})},{d_{\nu}({P}_{*})}\right\rangle$ on the right hand side can be bounded as
@@
while $-\left\langle{D(\pi_{*};\pi_{T})},{d_{\nu}({P}_{*})}\right\rangle$ can be dropped due to the positivity of Bregman divergences yielding
$\begin{split}\sum_{t=0}^{T-1}J({P}_{*})-J({P}_{t})\leq&\frac{2}{(1-\gamma)^{2}}\left(\bigl{\|}{r}\bigr{\|}_{\infty}+\sum_{t=0}^{T-1}\varepsilon_{t}\right)+\frac{1}{\eta(1-\gamma)}\left\langle{D(\pi_{*};\pi_{0})},{d_{\nu}({P}_{*})}\right\rangle.\end{split}$ (B.10)
@@
$J({P}_{*})-J({P}_{T})\leq\varepsilon_{T}+\frac{1}{T}\sum_{t=0}^{T-1}J({P}_{*})-J({P}_{t})+\varepsilon_{t}.$
@@
$J({P}_{*})-J({P}_{T})\leq\varepsilon_{T}+\frac{1}{T}\left[\left(1+\frac{2}{(1-\gamma)^{2}}\right)\sum_{t=0}^{T-1}\varepsilon_{t}+\frac{2\bigl{\|}{r}\bigr{\|}_{\infty}}{(1-\gamma)^{2}}+\frac{1}{\eta(1-\gamma)}\left\langle{D(\pi_{*};\pi_{0})},{d_{\nu}({P}_{*})}\right\rangle\right],$
@@ Appendix C POWR Convergence Rates
In this section we prove the convergence of POWR . To do so, we need to first show that under the choice of spaces ${\mathcal{F}}$ and ${\mathcal{G}}$ proposed in this work, the resulting PMD iterations are well defined. Then, we need to bound the approximation error of the estimates for the action-value functions of the iterates produced by the inexact PDM algorithm, which appear in the rates of Thm. 7.
@@ C.1 POWR is Well-defined
In order to guarantee that the iterations of POWR generate policies $\pi_{t}$ for which we can compute an estimator according to the formula in Prop. 3, we need to guarantee that all such policies are $({\mathcal{G}},{\mathcal{F}})$-compatible. In particular we restrict to the case of the separable spaces introduced in Prop. 4, for which it turns out that it is sufficient to show that all policies belong to the space ${\mathcal{H}}$ characterizing ${\mathcal{F}}={\mathcal{H}}\otimes{\mathcal{H}}$ and ${\mathcal{G}}=\mathbb{R}^{|\mathcal{A}|}\otimes{\mathcal{H}}$. The following results provide a candidate for choosing such space.
@@
We recall that Sobolev spaces [33] over a compact subset $\mathcal{X}$ of $\mathbb{R}^{D}$ are closed with respect to the operations of sum, multiplication, exponentiation or inversion (if the function is supported on the entire domain $\mathcal{X}$), namely for any two $f,f^{\prime}\in{\mathcal{H}}$, $f+f^{\prime},ff^{\prime},e^{f}\in{\mathcal{H}}$ and, if $f(x)>0$ for all $x\in{\mathcal{X}}$, $1/f\in{\mathcal{H}}$. This follows from by applying the chain rule and the boundedness of derivatives over the compact ${\mathcal{X}}$ (see for instance [42, Lemma E.2.2]). The proof follows by observing that the one step update $\pi_{t+1}$ in 13 is expressed precisely in terms of these operations and the hypothesis that $\pi_{t}(a|\cdot)$ and $\hat{q}_{t}(\cdot,a)$ belong to ${\mathcal{H}}$ for any $a\in\mathcal{A}$. ∎
@@
We proceed by induction. Since $\bar{q}(\cdot,a)\in{\mathcal{H}}$ we can apply the same reasoning in Thm. 5 to guarantee that $\pi_{0}(a|\cdot)\in{\mathcal{H}}$ for any $a\in\mathcal{A}$. Moreover, $\pi_{0}(a|\cot)>0$ for any $a\in\mathcal{A}$ since it is the (normalized) exponential of a function. Hence $\pi_{0}$ is $({\mathcal{G}},{\mathcal{F}})$-compatible. Therefore, $\hat{q}_{0}$ obtained according to Prop. 3 is well defined and belongs to ${\mathcal{G}}$, implying $\hat{q}_{0}(\cdot,a)\in{\mathcal{H}}$ for any $a\in\mathcal{A}$. Now, assume by inductive hypothesis that the policy $\pi_{t}(a|\cdot)$ generated by POWR at time $t$ and the corresponding estimator $\hat{q}_{t}(\cdot,a)$ of the action value function belong to ${\mathcal{H}}$ and that $\pi_{t}(a|x)>0$ for any $(x,a)\in\Omega$. Then, by Thm. 5 we have that also $\pi_{t+1}$ the solution to the PMD update in 13 belongs to ${\mathcal{H}}$ (and is therefore $({\mathcal{G}},{\mathcal{F}})$-compatible). Additionally, since $\pi_{t+1}$ can be expressed as the softmax of a (finite) sum of functions in ${\mathcal{H}}$, we have also $\pi_{t+1}(a|x)>0$ for al $(x,a)\in\Omega$, proving the inductive hypothesis and concluding the proof. ∎
The above corollary guarantees us that if we are able to learn our estimates for the action-value function in ${\mathcal{H}}$ a suitably regular Sobolev space, then POWR is well-defined. This is a necessary condition to then being able to study it’s theoretical behavior in our main result.
@@ C.2 Controlling the Action-value Estimation Error
We now show how to control the eximation error for the action-value funciton. we start by considering the following application of the (generalized) Simulation lemma in Corollary A.8.
@@
Recall that in the notation of these appendices the action value of a policy and its estimator via the world model CME framework are denoted $q_{\pi}=q({P})$ and $\hat{q}_{\pi}=q_{n}(\pi)$ respectively. We can apply Corollary A.8 to obtain
$q_{n}({P})-q({P})=({Id}-\gamma{T}_{n}{P})^{-1}(r_{n}-r)+\gamma({Id}-\gamma{T}_{n}{P})^{-1}({T}-{T}_{n})v({P}).$
@@
$\bigl{\|}{q_{n}({P})-q({P})}\bigr{\|}_{\infty}\leq\frac{1}{1-\gamma^{\prime}}\Big{[}\bigl{\|}{r_{n}-r}\bigr{\|}_{\infty}+\gamma\bigl{\|}{({T}-{T}_{n})v({P})}\bigr{\|}_{\infty}\Big{]},$
where $v({P}):={P}({Id}-\gamma{T}{P})^{-1}r$ is the value function of the MDP, and we used that $\gamma\left\|{{T}_{n}}\right\|<\gamma^{\prime}$. Because of Asm. 2, and ${P}$ being $({\mathcal{G}},{\mathcal{F}})$-compatible, it holds that $r\in{\mathcal{G}}$, $v({P})\in{\mathcal{F}}$, while Prop. 3 implies $r_{n}\in{\mathcal{G}}$, and $({T}-{T}_{n})v({P})\in{\mathcal{G}}$ as well. Therefore, using the reproducing property
@@
where we assumed a bounded kernel $\left\langle{\psi(x,a)},{\psi(x,a)}\right\rangle\leq C_{\psi}$ for all $(x,a)\in\Omega$. Similarly, for the term depending on ${T}_{n}-{T}$ we have
$\begin{split}\bigl{\|}{({T}-{T}_{n})v({P})}\bigr{\|}_{\infty}&=\sup_{(x,a)\in\Omega}|[({T}|_{{\mathcal{F}}}-{T}_{n})v({P})](x,a)|\\ &=\sup_{(x,a)\in\Omega}|\left\langle{\psi(x,a)},{({T}|_{{\mathcal{F}}}-{T}_{n})v({P})}\right\rangle_{{\mathcal{G}}}|\\ &=\sup_{(x,a)\in\Omega}\Big{|}\text{\rm Tr}\left[(v({P})\otimes\psi(x,a))({T}|_{{\mathcal{F}}}-{T}_{n})\right]\Big{|}\\ &\leq\bigl{\|}{{T}|_{{\mathcal{F}}}-{T}_{n}}\bigr{\|}_{{HS}}\sup_{(x,a)\in\Omega}|v({P})(x,a)|\\ &\leq\frac{\left\|{r}\right\|_{\infty}}{1-\gamma}\bigl{\|}{{T}|_{{\mathcal{F}}}-{T}_{n}}\bigr{\|}_{{HS}}.\end{split}$
@@
$\bigl{\|}{q_{n}({P})-q({P})}\bigr{\|}_{\infty}\leq\frac{1}{1-\gamma^{\prime}}\left[C_{\psi}\bigl{\|}{r_{n}-r}\bigr{\|}_{{\mathcal{G}}}+\frac{\gamma\bigl{\|}{r}\bigr{\|}_{\infty}}{1-\gamma}\bigl{\|}{{T}|_{{\mathcal{F}}}-{T}_{n}}\bigr{\|}_{{HS}}\right],$
@@
According to the result above, we can control the approximation error for the action value function in terms of the approximation errors $\left\|{r_{n}-r}\right\|_{{\mathcal{G}}}$ and $\bigl{\|}{{T}|_{{\mathcal{F}}}-{T}_{n}}\bigr{\|}_{{HS}}$. This can be done by leveraging state-of-the-art statistical learning rates for the ridge regression and CME estimators from [35, 18, 43]. The following lemma connects Asm. 2 with the notation used in [35] which enables us to use the required result.
@@
Let ${T}_{n}$ and $r_{n}$ the empirical estimators of the transfer operator ${T}$ and reward function $r$ as defined in Prop. 3, respectively. When ${P}$ is a $({\mathcal{G}},{\mathcal{F}})$-compatible policy as in Definition 1 and the strong source condition Asm. 2 is attained with parameter $\beta$, it holds
$\bigl{\|}{q_{n}({P})-q({P})}\bigr{\|}_{\infty}\leq O(\delta^{2}n^{-\alpha}),$ (C.2)
@@
The rate $\alpha_{r}\in\left(\frac{\beta}{2+2\beta},\frac{\beta}{1+2\beta}\right)$ is determined by the properties of the inclusion ${\mathcal{G}}\hookrightarrow B_{b}(\Omega)$, and the constant $c_{r}>0$ is independent of $n$ and $\delta$. Similarly, point (2.) of [18, Theorem 2] shows that under Asm. 2
$\bigl{\|}{{T}|_{{\mathcal{F}}}-{T}_{n}}\bigr{\|}_{{HS}({\mathcal{F}},{\mathcal{G}})}\leq\delta^{2}c_{{T}}n^{-\alpha_{{T}}}$ (C.4)
again with probability not less than $1-4e^{-\delta}$, rates $\alpha_{{T}}\in\left(\frac{\beta}{2+2\beta},\frac{\beta}{1+2\beta}\right)$ and with $c_{{T}}>0$ independent of $n$ and $\delta$. Combining every bound and denoting $\alpha:=\min(\alpha_{r},\alpha_{{T}})$, we conclude
$\bigl{\|}{q_{n}({P})-q({P})}\bigr{\|}_{\infty}\leq\frac{\delta^{2}}{1-\gamma^{\prime}}\left[C_{\psi}c_{r}+\frac{\gamma c_{{T}}\bigl{\|}{r}\bigr{\|}_{\infty}}{1-\gamma}\right]n^{-\alpha}=O(\delta^{2}n^{-\alpha}).$ (C.5)

### v2

@@ B.1 Convergence rates of PMD
As usual, in this proof we denote the estimated and exact action-value functions as $q_{n}(\mathsf{P}_{t}):=\hat{q}_{\pi_{t}}=(\mathsf{Id}-\gamma\mathsf{T}_{n}\mathsf{P}_{t})^{-1}r_{n}$ and $q(\mathsf{P}_{t}):=q_{\pi_{t}}=(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P}_{t})^{-1}r$, respectively. From hypothesis, Alg. 1 is well-defined since all policies it generates are $({\mathcal{G}},{\mathcal{F}})$-compatible. The resulting sequence of policies $(\pi_{t})_{t\in{\mathbb{N}}}$ are generated via the update rule 14 on the inexact action-value functions $q_{n}(\mathsf{P}_{t})$, as defined in 12. As the update rule 14 is a (measurable) minimizer of B.1 when $D$ equals the Kullback-Leibler divergence, the three-points Lemma B.2 with $\pi(x)=\pi_{*}(x)$ yields
$\left[(\mathsf{P}_{*}-\mathsf{P}_{t+1})q_{n}(\mathsf{P}_{t})\right](x)\leq\frac{1}{\eta}D(\pi_{*}(x);\pi_{t}(x))-\frac{1}{\eta}D(\pi_{*}(x);\pi_{t+1}(x)).$
Adding and subtracting the term $\left[(\mathsf{P}_{*}-\mathsf{P}_{t+1})q(\mathsf{P}_{t})\right](x)$, and bounding the remaining difference as $\left[(\mathsf{P}_{*}-\mathsf{P}_{t+1})(q(\mathsf{P}_{t})-q_{n}(\mathsf{P}_{t}))\right](x)\leq 2\varepsilon_{t}$ – see the derivation of B.8 – one gets
$\left[(\mathsf{P}_{*}-\mathsf{P}_{t+1})q(\mathsf{P}_{t})\right](x)\leq 2\varepsilon_{t}+\frac{1}{\eta}D(\pi_{*}(x);\pi_{t}(x))-\frac{1}{\eta}D(\pi_{*}(x);\pi_{t+1}(x)).$
Adding and subtracting $\mathsf{P}_{t}q(\mathsf{P}_{t})$ on the left side gives
$\left[(\mathsf{P}_{*}-\mathsf{P}_{t})q(\mathsf{P}_{t})\right](x)\leq\left[(\mathsf{P}_{t+1}-\mathsf{P}_{t})q(\mathsf{P}_{t})\right](x)+2\varepsilon_{t}+\frac{1}{\eta}D(\pi_{*}(x);\pi_{t}(x))-\frac{1}{\eta}D(\pi_{*}(x);\pi_{t+1}(x)),$
and integrating with respect to the positive measure $(\mathsf{Id}-\gamma\mathsf{P}_{*}\mathsf{T})^{-*}\nu$ and using the performance difference Lemma A.3 on the left hand side one has
$\begin{split}J(\mathsf{P}_{*})-J(\mathsf{P}_{t})\leq&\frac{1}{1-\gamma}\left\langle{(\mathsf{P}_{t+1}-\mathsf{P}_{t})q(\mathsf{P}_{t})+2\varepsilon_{t}},{d_{\nu}(\mathsf{P}_{*})}\right\rangle\\ &\frac{1}{\eta(1-\gamma)}\left\langle{D(\pi_{*};\pi_{t})-D(\pi_{*};\pi_{t+1})},{d_{\nu}(\mathsf{P}_{*})}\right\rangle,\end{split}$ (B.9)
where we used A.8 on the right-hand-side terms. Since $(\mathsf{P}_{t+1}-\mathsf{P}_{t})q(\mathsf{P}_{t})+2\varepsilon_{t}\geq 0$ because of B.7, we can use fact (1) from Lemma A.5 with $(\mathsf{Id}-\gamma\mathsf{P}_{t+1}\mathsf{T})^{-1}$ and the performance difference Lemma A.3 to get
$\begin{split}\left\langle{(\mathsf{P}_{t+1}-\mathsf{P}_{t})q(\mathsf{P}_{t})+2\varepsilon_{t}},{d_{\nu}(\mathsf{P}_{*})}\right\rangle&\leq\left\langle{(\mathsf{Id}-\gamma\mathsf{P}_{t+1}\mathsf{T})^{-1}\left[(\mathsf{P}_{t+1}-\mathsf{P}_{t})q(\mathsf{P}_{t})+2\varepsilon_{t}\right]},{d_{\nu}(\mathsf{P}_{*})}\right\rangle\\ &=\left\langle{\mathsf{P}_{t+1}q(\mathsf{P}_{t+1})},{d_{\nu}(\mathsf{P}_{*})}\right\rangle-\left\langle{\mathsf{P}_{t}q(\mathsf{P}_{t})},{d_{\nu}(\mathsf{P}_{*})}\right\rangle+\frac{2\varepsilon_{t}}{1-\gamma}.\\ \end{split}$
@@
$\begin{split}\sum_{t=0}^{T-1}J(\mathsf{P}_{*})-J(\mathsf{P}_{t})\leq&\frac{1}{1-\gamma}\left(\left\langle{\mathsf{P}_{T}q(\mathsf{P}_{T})},{d_{\nu}(\mathsf{P}_{*})}\right\rangle-\left\langle{\mathsf{P}_{0}q(\mathsf{P}_{0})},{d_{\nu}(\mathsf{P}_{*})}\right\rangle\right)+\frac{2}{(1-\gamma)^{2}}\sum_{t=0}^{T-1}\varepsilon_{t}\\ &\frac{1}{\eta(1-\gamma)}\left\langle{D(\pi_{*};\pi_{0})-D(\pi_{*};\pi_{T})},{d_{\nu}(\mathsf{P}_{*})}\right\rangle.\end{split}$
Using facts (3) and (4) from Lemma A.5 we have that the terms $\left\langle{\mathsf{P}q(\mathsf{P})},{d_{\nu}(\mathsf{P}_{*})}\right\rangle$ on the right hand side can be bounded as
@@
while $-\left\langle{D(\pi_{*};\pi_{T})},{d_{\nu}(\mathsf{P}_{*})}\right\rangle$ can be dropped due to the positivity of Bregman divergences yielding
$\begin{split}\sum_{t=0}^{T-1}J(\mathsf{P}_{*})-J(\mathsf{P}_{t})\leq&\frac{2}{(1-\gamma)^{2}}\left(\bigl{\|}{r}\bigr{\|}_{\infty}+\sum_{t=0}^{T-1}\varepsilon_{t}\right)+\frac{1}{\eta(1-\gamma)}\left\langle{D(\pi_{*};\pi_{0})},{d_{\nu}(\mathsf{P}_{*})}\right\rangle.\end{split}$ (B.10)
@@
$J(\mathsf{P}_{*})-J(\mathsf{P}_{T})\leq\varepsilon_{T}+\frac{1}{T}\sum_{t=0}^{T-1}J(\mathsf{P}_{*})-J(\mathsf{P}_{t})+\varepsilon_{t}.$
@@
$J(\mathsf{P}_{*})-J(\mathsf{P}_{T})\leq\varepsilon_{T}+\frac{1}{T}\left[\left(1+\frac{2}{(1-\gamma)^{2}}\right)\sum_{t=0}^{T-1}\varepsilon_{t}+\frac{2\bigl{\|}{r}\bigr{\|}_{\infty}}{(1-\gamma)^{2}}+\frac{1}{\eta(1-\gamma)}\left\langle{D(\pi_{*};\pi_{0})},{d_{\nu}(\mathsf{P}_{*})}\right\rangle\right],$
@@ Appendix C POWR Convergence Rates
In this section, we prove the convergence of POWR. To do so, we need to first show that under the choice of spaces ${\mathcal{F}}$ and ${\mathcal{G}}$ proposed in this work, the resulting PMD iterations are well defined. Then, we need to bound the approximation error of the estimates for the action-value functions of the iterates produced by the inexact PDM algorithm, which appear in the rates of Thm. 7.
@@ C.1 POWR  is Well-defined
In order to guarantee that the iterations of POWR  generate policies $\pi_{t}$ for which we can compute an estimator according to the formula in Prop. 3, we need to guarantee that all such policies are $({\mathcal{G}},{\mathcal{F}})$-compatible. In particular, we restrict to the case of the separable spaces introduced in Prop. 4, for which it turns out that it is sufficient to show that all policies belong to the space ${\mathcal{H}}$ characterizing ${\mathcal{F}}={\mathcal{H}}\otimes{\mathcal{H}}$ and ${\mathcal{G}}={\mathbb{R}}^{|\mathcal{A}|}\otimes{\mathcal{H}}$. The following results provide a candidate for choosing such a space.
@@
We recall that Sobolev spaces [37] over a compact subset $\mathcal{X}$ of ${\mathbb{R}}^{D}$ are closed with respect to the operations of sum, multiplication, exponentiation or inversion (if the function is supported on the entire domain $\mathcal{X}$), namely for any two $f,f^{\prime}\in{\mathcal{H}}$, $f+f^{\prime},ff^{\prime},e^{f}\in{\mathcal{H}}$ and, if $f(x)>0$ for all $x\in{\mathcal{X}}$, $1/f\in{\mathcal{H}}$. This follows by applying the chain rule and the boundedness of derivatives over the compact ${\mathcal{X}}$ (see for instance [45, Lemma E.2.2]). The proof follows by observing that the one-step update $\pi_{t+1}$ in 14 is expressed precisely in terms of these operations and the hypothesis that $\pi_{t}(a|\cdot)$ and $\hat{q}_{\pi_{t}}(\cdot,a)$ belong to ${\mathcal{H}}$ for any $a\in\mathcal{A}$. ∎
@@
We proceed by induction. Since $\bar{q}(\cdot,a)\in{\mathcal{H}}$ we can apply the same reasoning in Thm. 5 to guarantee that $\pi_{0}(a|\cdot)\in{\mathcal{H}}$ for any $a\in\mathcal{A}$. Moreover, $\pi_{0}(a|\cot)>0$ for any $a\in\mathcal{A}$ since it is the (normalized) exponential of a function. Hence $\pi_{0}$ is $({\mathcal{G}},{\mathcal{F}})$-compatible. Therefore, $\hat{q}_{0}$ obtained according to Prop. 3 is well defined and belongs to ${\mathcal{G}}$, implying $\hat{q}_{0}(\cdot,a)\in{\mathcal{H}}$ for any $a\in\mathcal{A}$. Now, assume by the inductive hypothesis that the policy $\pi_{t}(a|\cdot)$ generated by POWR at time $t$ and the corresponding estimator $\hat{q}_{\pi_{t}}(\cdot,a)$ of the action value function belong to ${\mathcal{H}}$ and that $\pi_{t}(a|x)>0$ for any $(x,a)\in\Omega$. Then, by Thm. 5 we have that also $\pi_{t+1}$ the solution to the PMD update in 14 belongs to ${\mathcal{H}}$ (and is therefore $({\mathcal{G}},{\mathcal{F}})$-compatible). Additionally, since $\pi_{t+1}$ can be expressed as the softmax of a (finite) sum of functions in ${\mathcal{H}}$, we have also $\pi_{t+1}(a|x)>0$ for al $(x,a)\in\Omega$, proving the inductive hypothesis and concluding the proof. ∎
The above corollary guarantees us that if we are able to learn our estimates for the action-value function in ${\mathcal{H}}$ a suitably regular Sobolev space, then POWR  is well-defined. This is a necessary condition to then being able to study its theoretical behavior in our main result.
@@ C.2 Controlling the Action-value Estimation Error
We now show how to control the estimation error for the action-value function. We start by considering the following application of the (generalized) Simulation lemma in Corollary A.7.
@@
Recall that in the notation of these appendices, the action value of a policy and its estimator via the world model CME framework are denoted $q_{\pi}=q(\mathsf{P})$ and $\hat{q}_{\pi}=q_{n}(\mathsf{P})$ respectively. We can apply Corollary A.7 to obtain
$q_{n}(\mathsf{P})-q(\mathsf{P})=(\mathsf{Id}-\gamma\mathsf{T}_{n}\mathsf{P})^{-1}(r_{n}-r)+\gamma(\mathsf{Id}-\gamma\mathsf{T}_{n}\mathsf{P})^{-1}(\mathsf{T}-\mathsf{T}_{n})v(\mathsf{P}).$
@@
$\bigl{\|}{q_{n}(\mathsf{P})-q(\mathsf{P})}\bigr{\|}_{\infty}\leq\frac{1}{1-\gamma^{\prime}}\Big{[}\bigl{\|}{r_{n}-r}\bigr{\|}_{\infty}+\gamma\bigl{\|}{(\mathsf{T}-\mathsf{T}_{n})v(\mathsf{P})}\bigr{\|}_{\infty}\Big{]},$
where $v(\mathsf{P}):=\mathsf{P}(\mathsf{Id}-\gamma\mathsf{T}\mathsf{P})^{-1}r$ is the value function of the MDP, and we used that $\gamma\left\|{\mathsf{T}_{n}}\right\|<\gamma^{\prime}$. Because of Asm. 1, $r\in{\mathcal{G}}$, and $\mathsf{P}$ being $({\mathcal{G}},{\mathcal{F}})$-compatible, it holds that $v(\mathsf{P})\in{\mathcal{F}}$, while Prop. 3 implies $r_{n}\in{\mathcal{G}}$, and $(\mathsf{T}-\mathsf{T}_{n})v(\mathsf{P})\in{\mathcal{G}}$ as well. Therefore, using the reproducing property
@@
where we assumed a bounded kernel $\left\langle{\psi(x,a)},{\psi(x,a)}\right\rangle\leq C_{\psi}$ for all $(x,a)\in\Omega$. Similarly, for the term depending on $\mathsf{T}_{n}-\mathsf{T}$ we have
$\begin{split}\bigl{\|}{(\mathsf{T}-\mathsf{T}_{n})v(\mathsf{P})}\bigr{\|}_{\infty}&=\sup_{(x,a)\in\Omega}|[(\mathsf{T}|_{{\mathcal{F}}}-\mathsf{T}_{n})v(\mathsf{P})](x,a)|\\ &=\sup_{(x,a)\in\Omega}|\left\langle{\psi(x,a)},{(\mathsf{T}|_{{\mathcal{F}}}-\mathsf{T}_{n})v(\mathsf{P})}\right\rangle_{{\mathcal{G}}}|\\ &=\sup_{(x,a)\in\Omega}\Big{|}\textrm{Tr}\left[(v(\mathsf{P})\otimes\psi(x,a))(\mathsf{T}|_{{\mathcal{F}}}-\mathsf{T}_{n})\right]\Big{|}\\ &\leq\bigl{\|}{\mathsf{T}|_{{\mathcal{F}}}-\mathsf{T}_{n}}\bigr{\|}_{\mathsf{HS}}\sup_{(x,a)\in\Omega}|v(\mathsf{P})(x,a)|\\ &\leq\frac{\left\|{r}\right\|_{\infty}}{1-\gamma}\bigl{\|}{\mathsf{T}|_{{\mathcal{F}}}-\mathsf{T}_{n}}\bigr{\|}_{\mathsf{HS}}.\end{split}$
@@
$\bigl{\|}{q_{n}(\mathsf{P})-q(\mathsf{P})}\bigr{\|}_{\infty}\leq\frac{1}{1-\gamma^{\prime}}\left[C_{\psi}\bigl{\|}{r_{n}-r}\bigr{\|}_{{\mathcal{G}}}+\frac{\gamma\bigl{\|}{r}\bigr{\|}_{\infty}}{1-\gamma}\bigl{\|}{\mathsf{T}|_{{\mathcal{F}}}-\mathsf{T}_{n}}\bigr{\|}_{\mathsf{HS}}\right],$
@@
According to the result above, we can control the approximation error for the action value function in terms of the approximation errors $\left\|{r_{n}-r}\right\|_{{\mathcal{G}}}$ and $\bigl{\|}{\mathsf{T}|_{{\mathcal{F}}}-\mathsf{T}_{n}}\bigr{\|}_{\mathsf{HS}}$. This can be done by leveraging state-of-the-art statistical learning rates for the ridge regression and CME estimators from [39, 21, 46]. The following lemma connects Asm. 2 with the notation used in [39] which enables us to use the required result.
@@
Let $\mathsf{T}_{n}$ and $r_{n}$ the empirical estimators of the transfer operator $\mathsf{T}$ and reward function $r$ as defined in Prop. 3, respectively. When $\mathsf{P}$ is a $({\mathcal{G}},{\mathcal{F}})$-compatible policy as in Def. 1 and the strong source condition Asm. 2 is attained with parameter $\beta$, it holds
$\bigl{\|}{q_{n}(\mathsf{P})-q(\mathsf{P})}\bigr{\|}_{\infty}\leq O(\delta^{2}n^{-\alpha}),$ (C.2)
@@
The rate $\alpha_{r}\in\left(\frac{\beta}{2+2\beta},\frac{\beta}{1+2\beta}\right)$ is determined by the properties of the inclusion ${\mathcal{G}}\hookrightarrow B_{b}(\Omega)$, and the constant $c_{r}>0$ is independent of $n$ and $\delta$. Similarly, point (2.) of [21, Theorem 2] shows that under Asm. 2
$\bigl{\|}{\mathsf{T}|_{{\mathcal{F}}}-\mathsf{T}_{n}}\bigr{\|}_{\mathsf{HS}({\mathcal{F}},{\mathcal{G}})}\leq\delta^{2}c_{\mathsf{T}}n^{-\alpha_{\mathsf{T}}}$ (C.4)
again with probability not less than $1-4e^{-\delta}$, rates $\alpha_{\mathsf{T}}\in\left(\frac{\beta}{2+2\beta},\frac{\beta}{1+2\beta}\right)$ and with $c_{\mathsf{T}}>0$ independent of $n$ and $\delta$. Combining every bound and denoting $\alpha:=\min(\alpha_{r},\alpha_{\mathsf{T}})$, we conclude
$\bigl{\|}{q_{n}(\mathsf{P})-q(\mathsf{P})}\bigr{\|}_{\infty}\leq\frac{\delta^{2}}{1-\gamma^{\prime}}\left[C_{\psi}c_{r}+\frac{\gamma c_{\mathsf{T}}\bigl{\|}{r}\bigr{\|}_{\infty}}{1-\gamma}\right]n^{-\alpha}=O(\delta^{2}n^{-\alpha}).$ (C.5)